OpenStudy (anonymous):

Why any number raise to power o equal 1?

OpenStudy (shadowfiend):

That's sort of a tough question. It's basically by definition.

OpenStudy (anonymous):

its by exponents

OpenStudy (shadowfiend):

If you want to think about it, raising a number A to the power B means `number A, B times, with all those times multiplied together'. If you have it 0 times, you don't have it, so we say it's one.

OpenStudy (anonymous):

A number raised to the zero power is always one because.. it's just how it works.. Look at this: 1= 3^0/3^0, can we agree on that? A number divided by itself is 1. However, for that to be true, we know that if 3^0 is 0, we have 0/0, a mathematical impossibility.

OpenStudy (anonymous):

Additionally, exponants are a sequence. 2^1 = 2 2^2 = 2*2 2^3 = 2*2*2 If we follow the rule, 2^0 has to be 1 in order to follow the sequence. I refer you to this page: http://mathforum.org/dr.math/faq/faq.number.to.0power.html

OpenStudy (anonymous):

THANK YOU FOR THE HELP

OpenStudy (shadowfiend):

No problem, glad to help :)